# Monte Carlo Physics-informed neural networks for multiscale heat conduction via phonon Boltzmann transport equation

### v3

@@ 3.3 Quasi-2D square heat conduction
As mentioned, the method in [17] is one of the most computationally and memory-efficient approaches for solving stationary BTE. We now perform a comparison of the computational and the memory cost between the MC-PINNs and the numerical method [17] that is used to generate the reference solution here. Specifically, we compare the computational time and the memory usage using two representative cases, i.e., Kn $=0.1$ and 10.
We first discuss the computational time for MC-PINNs. As illustrated in Table 5, the computational time of MC-PINNs (second column in Table 5) is about 7 and 1.3 times that in the numerical method [17] for Kn $=0.1$ and 10, respectively. In the deep learning community, the transfer learning technique is a widely used approach to speed up the training of DNNs. Here we can also use it to speedup the training of MC-PINNs. Specifically, we first train the MC-PINNs from scratch for Kn $=1$, and then save the hyperparameters, i.e., weights and biases, in order to reuse them in other cases. In particular, we initialize the MC-PINNs for Kn $=0.1$ and 10 using the hyperparameters from the pre-trained MC-PINNs for Kn $=1$ and further train them for an additional number of steps until the convergence is reached. The relative $L_{2}$ errors for these two cases with transfer learning are $2.40\%$ and $2.6\%$, which are similar to the cases without transfer learning. Also, as we can see in Table 5, we can achieve about 3-6 times speedup with transfer learning, and the computational times for these two cases are now comparable to the numerical method in [17].
@@ 4 Summary
In this study, we develop a Monte Carlo physics-informed neural networks (MC-PINNs) for solving the phonon Boltzmann transport equation (BTE) to model the multiscale heat conduction in solid materials. Specifically, we utilize a deep neural network to approximate the energy density function in BTE and encode the BTE as well as the boundary/initial conditions via the automatic differentiation, in MC-PINNs. In addition, we propose a novel two-step sampling approach to address the issues like “curse of dimensionality” and the inaccurate estimate of the equilibrium distribution functions in the existing sampling strategies that are widely used in PINNs. In particular, we first randomly sample a certain number of points in the temporal-spatial domains (Step I), and then we draw another number of random points in the solid angular domain (Step II). The residual points in MC-PINNs are constructed using the tensor product given the points generated at the above two steps. We conduct a series of numerical experiments to validate the accuracy of the MC-PINNs, including the steady quasi-1D/quasi-2D/3D as well as unsteady quasi-1D heat conduction problems spanning the diffusive and ballistic regimes. The results show the MC-PINNs is capable of achieving good accuracy for the multiscale heat conduction from 1D to 3D. Furthermore, the present method requires quite a small number of discrete points in the solid angular domain than in the deterministic methods for solving the BTE, especially for scenarios in the ballistic regimes. For instance, we only employ 64 points in the solid angular domain at each training step in MC-PINNs for Kn ranging from 0.1 to 10 in the 3D case. In contrast, we have $80\times 80$ discrete points in the solid angular domain of the numerical method to obtain converged results. As a result, the memory cost in the MC-PINN is about 1/6 of that in the numerical method. Further, we are able to achieve a similar computational efficiency for MC-PINNs to the numerical method in the 3D test case as we utilize the transfer learning technique in the MC-PINNs.
@@
We would like to discuss that the MC-PINN developed in this work is mesh-free in the temporal-spatial-angular domain, which is promising in real-world applications with complicated domains, for instance, the design of microdevices. In addition, the MC-PINN is capable of providing good accuracy for heat conduction ranging from diffusive to ballistic with only a small number of random discrete points in the solid angular domain at each training step. It can save lots of effort in numerical methods since the optimal discretization of the solid angular space at different regimes is generally different and should be carefully designed to ensure convergence in these methods. Also, the memory efficacy and the flexibility of treatment for the solid angular space in the present method make it a promising tool for modeling multiscale heat conduction in large-scale problems. Finally, we note that the MC-PINNs with the two-step sampling approach can be seamlessly extended to solve kinetic equations used in other disciplines, e.g., the Boltzmann-BGK model for multiscale flows, the radiation transfer equation for multiscale heat transfer, etc. We will leave these interesting topics as future studies.
Acknowledgement
@@
QL, XM, and ZG acknowledge the support of the National Natural Science Foundation of China (No. 12201229) and the Interdisciplinary Research Program of HUST (No. 2023JCYJ002 and 2024JCYJ003). CZ was supported by the National Natural Science Foundation of China (No. 12147122). XM would also like to acknowledge the support of Xiaomi Young Talents Program. The authors acknowledge Beijng PARATERA Tech CO., Ltd. for the HPC resources.
Appendix A Details on the computations
In all the test cases of Sec. 3, the Adam optimizer with a learning rate $10^{-3}$ is employed for the training of PINNs. More details on the architectures of MC-PINNs, etc., are present in Table 7.
Table 7: Architecture and training step for MC-PINNs in each case.
@@ Appendix A Details on the computations
Appendix A $20\times 4$ tanh 50,000
@@
Table 8: Grid size used in the numerical method for generating reference solutions in each case. For the 3D case in Sec. 3.4, we use $40\times 40$ and $80\times 80$ points in the solid angular domain for $\rm{Kn}=0.1/1$ and 10, respectively.
@@
Appendix A - 40 100
Appendix B Ablation study
@@
Figure 15: Quasi-1D film heat conduction: Relative $L_{2}$ errors of the temperature with different numbers of points in the solid angular space at each training step in PINNs. (a) $\rm{Kn}=1$, and (b) $\rm{Kn=10}$. Error bars in each figure represent the standard deviation calculated from 5 independent runs with different random seeds. The blue dashed line denotes the relative $L_{2}$ error for $1\%$.
We further present the loss histories for the representative case $\rm{Kn}=1.0$ with $B_{\bm{s}}=2$ and 16 in Fig. 16. As shown, the loss for MC-PINNs with 16 random points in the solid angular space after 50,000 training steps is about two orders smaller than that with 2 random points in the solid angular domain. Also, the fluctuations in the loss are smaller for the case with a larger number of points in the solid angular space at each training step.
Figure 16: Loss history for $B_{\bm{s}}=2$ and $B_{\bm{s}}=16$ at $\rm{Kn=1.0}$.

### v4

@@ 3.3 Quasi-2D square heat conduction
As mentioned, the method in [18] is one of the most efficient approaches for solving stationary BTE for all Knudsen numbers. We now perform a comparison of the computational and the memory cost between the MC-PINNs and the numerical method [18] that is used to generate the reference solution here. Specifically, we compare the computational time and the memory usage using two representative cases, i.e., Kn $=0.1$ and 10.
We first discuss the computational time for MC-PINNs. As illustrated in Table 5, the computational time of MC-PINNs trained from scratch (second column in Table 5) is about 7 and 1.3 times of that in the numerical method [18] for Kn $=0.1$ and 10, respectively. In the deep learning community, the transfer learning technique is a widely used approach to speed up the training of DNNs. Here we can also use it to speedup the training of MC-PINNs. Specifically, we first train the MC-PINNs from scratch for Kn $=1$, and then save the hyperparameters, i.e., weights and biases, in order to reuse them in other cases. In particular, we initialize the MC-PINNs for Kn $=0.1$ and 10 using the hyperparameters from the pre-trained MC-PINNs for Kn $=1$ and further train them for an additional number of steps until convergence. The relative $L_{2}$ errors for these two cases with transfer learning are $2.40\%$ and $2.60\%$, which are similar to the cases without transfer learning. Also, as we can see in Table 5, we can achieve about 3-6 times speedup with transfer learning, and the computational times for these two cases are now comparable to the numerical method in [18].
@@ 4 Summary
In this study, we develop Monte Carlo physics-informed neural networks (MC-PINNs) for solving the phonon Boltzmann transport equation (BTE) to model the multiscale heat conduction in solid materials. Specifically, we utilize a deep neural network to approximate the energy density function in BTE and encode the BTE as well as the boundary/initial conditions via the automatic differentiation, in MC-PINNs. In addition, we propose a novel two-step sampling approach to address the issues like “curse of dimensionality” and the inaccurate estimate of the equilibrium distribution functions in the existing sampling strategies that are widely used in PINNs. In particular, we first randomly sample a certain number of points in the temporal-spatial domains (Step I), and then we draw another number of random points in the solid angular domain (Step II). The residual points at each training step in MC-PINNs are constructed using the tensor product given the points generated at the above two steps. We conduct a series of numerical experiments to validate the accuracy of the MC-PINNs, including the steady quasi-1D/quasi-2D/3D as well as unsteady quasi-1D heat conduction problems spanning the diffusive and ballistic regimes. The results show the MC-PINNs is capable of achieving good accuracy for the multiscale heat conduction from 1D to 3D. Furthermore, the present method requires quite a small number of points in the solid angular domain than in the deterministic methods for solving the BTE, especially for scenarios in the ballistic regimes. For instance, we employ only 64 points in the solid angular domain at each training step in MC-PINNs for Kn ranging from 0.1 to 10 in the 3D case. In contrast, we have $80\times 80$ discrete points in the solid angular domain of the numerical method to obtain converged results. The memory cost in the MC-PINNs is about 1/6 of that in the numerical method. Further, we are able to achieve similar computational efficiency for MC-PINNs compared to the numerical method in the 3D test case as we utilize the transfer learning technique in the MC-PINNs.
@@
We would like to discuss that the MC-PINNs developed in this work is mesh-free in the temporal-spatial-angular domain, which is promising in real-world applications with complicated domains, for instance, the design of microdevices. In addition, the MC-PINNs is capable of providing good accuracy for heat conduction ranging from diffusive to ballistic with only a small number of random discrete points in the solid angular domain at each training step. It can save lots of effort in numerical methods since the optimal discretization of the solid angular space at different regimes is generally different and should be carefully designed to ensure convergence in these methods. Also, the memory efficacy and the flexibility of treatment for the solid angular space in the present method make it a promising tool for modeling multiscale heat conduction in large-scale problems. Finally, we note that the MC-PINNs with the two-step sampling approach can be seamlessly extended to solve kinetic equations used in other disciplines, e.g., the Boltzmann-BGK model for multiscale flows, the radiation transfer equation for multiscale heat transfer, etc. We will leave these interesting topics as future studies.
Acknowledgements
@@
Q. L., X. M., and Z. G. acknowledge the support of the National Natural Science Foundation of China (No. 12201229) and the Interdisciplinary Research Program of HUST (No. 2024JCYJ003 and 2023JCYJ002). C. Z. was supported by the National Natural Science Foundation of China (No. 12147122). X. M. would also like to acknowledge the support of Xiaomi Young Talents Program. The authors acknowledge the Beijng PARATERA Tech CO., Ltd. for the HPC resources.
Appendix A Gradients unbiasedness and convergence for the two-step sampling method in MC-PINNs
In this section, we follow [41] to conduct a theoretical analysis of the unbiasedness of the gradient and the convergence of the two-sampling method in MC-PINNs. Note that the sampling approach to compute the losses for the boundary and/or initial conditions are the same as the widely used minibatch training in the deep learning community, we thus only focus on the analysis of the residual loss in what follows.
We begin with the optimization trajectory of the gradient descent:
$\theta_{n+1}=\theta_{n}-\gamma\nabla f(\theta_{n}),$ (12)
where $\theta_{n}$ denotes the parameters at $n$-th training step, $\gamma$ is the learning rate, and $f(\theta_{n})$ is the loss function for the residual of the governing equation with $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$. Assuming that the full batch is constructed using sufficient points from the temporal-spatial and angular domains via a tensor product. The full-batch gradient can then be expressed as:
$\nabla f(\theta)=\frac{1}{N_{t,\bm{x}}N_{\bm{s}}}\sum_{i=1}^{N_{t,\bm{x}}}\sum_{j=1}^{N_{\bm{s}}}\nabla f_{i,j}(\theta),$ (13)
where $N_{t,\bm{x}}$ and $N_{\bm{s}}$ represent the numbers of full batch points in the temporal-spatial and angular domains, respectively.
In the two-step sampling approach, we randomly select two subsets from the full batches in the temporal-spatial and angular domains, respectively. The corresponding optimization trajectory for the two-step sampling method can be expressed as:
$\theta_{n+1}=\theta_{n}-\gamma\nabla f_{B_{t,\bm{x}},B_{\bm{s}}}(\theta_{n}),$ (14)
where
$\nabla f_{B_{t,\bm{x}},B_{\bm{s}}}(\theta)=\frac{1}{B_{t,\bm{x}}B_{\bm{s}}}\sum_{i=1}^{B_{t,\bm{x}}}\sum_{j=1}^{B_{\bm{s}}}\nabla f_{i,j}(\theta_{n}).$ (15)
Here, $B_{t,\bm{x}}$ and $B_{\bm{s}}$ denote the number of points in subsets in the temporal-spatial and angular space domains, respectively. The expectation of the gradient in Eq. (15) can be derived as follows:
$\mathbb{E}[\nabla f_{B_{t,\bm{x}},B_{\bm{s}}}(\theta)]$ $=\mathbb{E}\left[\frac{1}{B_{t,\bm{x}}B_{\bm{s}}}\sum_{i=1}^{B_{t,\bm{x}}}\sum_{j=1}^{B_{\bm{s}}}\nabla f_{i,j}(\theta)\right]$ (16)
$=\frac{1}{B_{t,\bm{x}}B_{\bm{s}}}\sum_{i=1}^{B_{t,\bm{x}}}\sum_{j=1}^{B_{\bm{s}}}\mathbb{E}[\nabla f_{i,j}(\theta)]$
$=\frac{1}{B_{t,\bm{x}}B_{\bm{s}}}\sum_{i=1}^{B_{t,\bm{x}}}\sum_{j=1}^{B_{\bm{s}}}\nabla f(\theta)$
$=\nabla f(\theta),$
which demonstrates that the gradient obtained by the two-step sampling method is unbiased.
We now conduct the analysis on the convergence of the two-step sampling method in MC-PINNs. Since $f$ is continuous and the number of full batch training points is finite, we can define the gradient noise in the two-step sampling method as follows according to [41] :
$\sigma^{*}=\inf_{\theta^{*}\in\mathop{\arg\min}f}\mathbb{V}[\nabla f_{B_{t,\bm{x}},B_{\bm{s}}}(\theta^{*})],$ (17)
where $\theta^{*}\in\mathop{\arg\min}f$ denotes the NN parameters when $f$ reaches the global minimum. Based on the trajectory in Eq. (14), we can obtain the following equation:
$\|\theta_{n+1}-\theta^{*}\|^{2}$ $=\|\theta_{n}-\gamma\nabla f_{B_{t,\bm{x}},B_{\bm{s}}}(\theta)-\theta^{*}\|^{2}$ (18)
$=\|\theta_{n}-\theta^{*}\|^{2}-2\gamma\langle\theta_{n}-\theta^{*},\nabla f_{B_{t,\bm{x}},B_{\bm{s}}}(\theta_{n})\rangle+\gamma^{2}\|\nabla f_{B_{t,\bm{x}},B_{\bm{s}}}(\theta_{n})\|^{2}.$
Taking the expectation conditioned on $\theta_{n}$ and utilizing the unbiasedness of $\nabla f_{B_{t,\bm{x}},B_{\bm{s}}}$ in Eq. (16), we obtain:
$\mathbb{E}[\|\theta_{n+1}-\theta^{*}\|^{2}|\theta_{n}]$ $=\|\theta_{n}-\theta^{*}\|^{2}-2\gamma\langle\theta_{n}-\theta^{*},\nabla f(\theta_{n})\rangle+\gamma^{2}\mathbb{E}[\|\nabla f_{B_{t,\bm{x}},B_{\bm{s}}}(\theta_{n})\|^{2}|\theta_{n}].$ (19)
Assuming that $f$ is a convex and differentiable function, we apply Lemma 2.8 in [41], which states:
$f(\theta^{*})-f(\theta_{n})\geq\langle\theta^{*}-\theta_{n},\nabla f(\theta_{n})\rangle,$ (20)
we can then obtain the following inequality:
$\mathbb{E}[\|\theta_{n+1}-\theta^{*}\|^{2}|\theta_{n}]\leq\|\theta_{n}-\theta^{*}\|^{2}-2\gamma(f(\theta_{n})-f(\theta^{*}))+\gamma^{2}\mathbb{E}[\|\nabla f_{B_{t,\bm{x}},B_{\bm{s}}}(\theta_{n})\|^{2}|\theta_{n}].$ (21)
We then assume that $f$ in the minibatch is $L_{b}$-smooth in expectation, which is defined for $\theta_{i},\theta_{j}\in\mathbb{R}^{d}$ as:
$\frac{1}{2L_{b}}\mathbb{E}[\|\nabla f_{B_{t,\bm{x}},B_{\bm{s}}}(\theta_{i})-\nabla f_{B_{t,\bm{x}},B_{\bm{s}}}(\theta_{j})\|^{2}]\leq f(\theta_{i})-f(\theta_{j})-\langle\nabla f(\theta_{j}),\theta_{i}-\theta_{j}),$ (22)
where $L_{b}$ is the Lipschitz constant. Then, following the Lemma 6.7 in [41], we can further obtain that:
$\mathbb{E}[\|\nabla f_{B_{t,\bm{x}},B_{\bm{s}}}(\theta)\|^{2}]\leq 4L_{b}(f(\theta)-f(\theta^{*}))+2\sigma^{*}.$ (23)
Consequently, the inequality Eq. (21) can be rewritten as:
$\mathbb{E}[\|\theta_{n+1}-\theta^{*}\|^{2}|\theta_{n}]\leq\|\theta_{n}-\theta^{*}\|^{2}+2\gamma(2\gamma L_{b}-1)(f(\theta_{n})-f(\theta^{*}))+2\gamma^{2}\sigma^{*}.$ (24)
We now reach the following inequality if we set the learning rate to satisfy $0<\gamma\leq 1/4L_{b}$ [41]:
$\mathbb{E}[\|\theta_{n+1}-\theta^{*}\|^{2}|\theta_{n}]\leq\|\theta_{n}-\theta^{*}\|^{2}-\gamma(f(\theta_{n})-f(\theta^{*}))+2\gamma^{2}\sigma^{*}.$ (25)
We then take expectation to obtain the following inequality:
$\mathbb{E}[\|\theta_{n+1}-\theta^{*}\|^{2}]\leq\mathbb{E}[\|\theta_{n}-\theta^{*}\|^{2}]-\gamma\mathbb{E}[f(\theta_{n})-f(\theta^{*})]+2\gamma^{2}\sigma^{*}.$ (26)
Summing over $i=0,1,...,n-1$, the above inequality can be written as:
$\mathbb{E}[\|\theta_{n}-\theta^{*}\|^{2}]\leq\|\theta_{0}-\theta^{*}\|^{2}-\gamma\sum_{i=0}^{n-1}\mathbb{E}[f(\theta_{i})-f(\theta^{*})]+2n\gamma^{2}\sigma^{*}.$ (27)
Since $\mathbb{E}[\|\theta_{n}-\theta^{*}\|^{2}]\geq 0$, we now obtain:
$\sum_{i=0}^{n-1}\mathbb{E}[f(\theta_{i})-f(\theta^{*})]\leq\frac{\|\theta_{0}-\theta^{*}\|^{2}}{\gamma}+2n\gamma\sigma^{*}.$ (28)
Defining $\bar{\theta}_{n}=\frac{1}{n}\sum_{i=0}^{n-1}\theta_{i}$ and invoking Jensen’s inequality for the convex function $f$ gives:
$\mathbb{E}[f(\bar{\theta}_{n})-f(\theta^{*})]\leq\frac{\|\theta_{0}-\theta^{*}\|^{2}}{\gamma n}+2\gamma\sigma^{*}.$ (29)
Specifically, when the training step $n$ is fixed and $n\geq 1$, we set $\gamma=\gamma_{0}/\sqrt{n}$:
$\mathbb{E}[f(\bar{\theta}_{n})-f(\theta^{*})]\leq\frac{\|\theta_{0}-\theta^{*}\|^{2}}{\gamma_{0}\sqrt{n}}+\frac{2\gamma_{0}\sigma^{*}}{\sqrt{n}}=\mathcal{O}(\frac{1}{\sqrt{n}}),$ (30)
where $\gamma_{0}$ is a constant with $0<\gamma_{0}\leq 1/4L_{b}$. The above results demonstrate the convergence of the two-sampling approach as the number of training steps $n$ becomes sufficiently large.
It is important to note that the optimization of PINNs generally involves a non-convex loss landscape. However, as reported in [42], when the PINNs converge to a local minimum at $\theta^{**}$, where the Hessian matrix of the loss function at $\theta^{**}$ is positive definite [43], the predicted accuracy is in general good. Further, the convergence to a good local minimum requires appropriate initialization of the neural networks. Empirically, the widely used Xavier’s initialization [44], which is also employed in this study, is found to be able to provide good initializations in MC-PINNs.
Appendix B Details on the computations
In all the test cases of Sec. 3, the Adam optimizer with a learning rate $10^{-3}$ is employed for the training of PINNs. More details on the architectures of MC-PINNs, etc., are present in Table 7. In addition, 64 points are utilized to compute the equilibrium in each case using the Monte Carlo integration. We note that we have tested the accuracy with more than 64 points, which show little difference from the results presented in this study. Therefore, we employ 64 points for computational efficiency.
Table 7: Architectures and training steps for MC-PINNs in each case.
@@ Appendix B Details on the computations
Appendix C $20\times 4$ tanh 50,000
@@
Table 8: Grid size used in the numerical method for generating reference solutions in each case. For the 3D case in Sec. 3.4, we use $40\times 40$ and $80\times 80$ points in the solid angular domain for $\rm{Kn}=0.1$, 1 and 10, respectively.
@@
Appendix B - 40 100
Appendix C Ablation study
@@
Figure 16: Quasi-1D film heat conduction: Relative $L_{2}$ errors of the temperature with different numbers of points in the solid angular space at each training step in MC-PINNs. (a) $\rm{Kn}=0.01$, (b) $\rm{Kn=1}$ and (c) $\rm{Kn=10}$. Error bars in each figure represent the standard deviation calculated from 5 independent runs with different random seeds.
We further present the loss histories for the representative case $\rm{Kn}=1.0$ with $B_{\bm{s}}=2$ and $16$ in Fig. 17. As shown, the loss for MC-PINNs with 16 random points in the solid angular space after 50,000 training steps is smaller than that with 2 random points in the solid angular domain. Also, the fluctuations in the loss are smaller for the case with a larger number of points in the solid angular space at each training step.
Figure 17: Loss history for $B_{\bm{s}}=2$ and $B_{\bm{s}}=16$ at $\rm{Kn=1}$.